1. (6 points) Find an optimal solution for the following transportation problem using the minimal cost method and the transportation algorithm: Minimize lahi + 2x12 + 2x13 + 4x21 + 3x22 + 4x23 + 4x31 + 1x32 + 3x33, subject to the constraints X11 + X12 + X13 = 100. x21 +x22 +x23 = 50. r31 + 232 +x33 100 x11 + 2'21 +2'3,-150. 12 22+32-50 x13 + x23 + x33-50. for all i, j = 1.2.3. xij > 0,
3. Consider the same transportation problem as in Question 3. The problem has an optimal solution given by 12 4,1 8, 1 8, 2 10,3 12,3 8 (a) (3 points) Solve the dual problem. Hint: Make use of complementary slackness to obtain a system of linear equations with one free variable. Then consider the equality of objective functions (b) (1 point) Which statement is true? The dual problem has □ a unique olt mal solution □ solutions □ finitely many optimal no optimal solution ם infinitely many optimal solutions
Answer #1

here the variables are denoted as follows:

x11 = x1                   x31 = x7

x12 = x2                  x32 = x8

x13 = x3                  x33 = x9

x21 = x4

x22 = x5

x23 = x6

Min Z = x1 + 2 x2 + 2 x3 + 4 x4 + 3 x5 + 4 x6 + 4 x7 + x8 + 3 x9
subject to
x1 + x2 + x3 = 100
x4 + x5 + x6 = 50
x7 + x8 + x9 = 100
x1 + x4 + x7 = 150
x2 + x5 + x8 = 50
x3 + x6 + x9 = 50
and x1,x2,x3,x4,x5,x6,x7,x8,x9≥0;



The problem is converted to canonical form by adding slack, surplus and artificial variables as appropriate

1. As the constraint-1 is of type '=' we should add artificial variable A1

2. As the constraint-2 is of type '=' we should add artificial variable A2

3. As the constraint-3 is of type '=' we should add artificial variable A3

4. As the constraint-4 is of type '=' we should add artificial variable A4

5. As the constraint-5 is of type '=' we should add artificial variable A5

6. As the constraint-6 is of type '=' we should add artificial variable A6

After introducing artificial variables

Min Z = x1 + 2 x2 + 2 x3 + 4 x4 + 3 x5 + 4 x6 + 4 x7 + x8 + 3 x9 + M A1 + M A2 + M A3 + M A4 + M A5 + M A6
subject to
x1 + x2 + x3 + A1 = 100
x4 + x5 + x6 + A2 = 50
x7 + x8 + x9 + A3 = 100
x1 + x4 + x7 + A4 = 150
x2 + x5 + x8 + A5 = 50
x3 + x6 + x9 + A6 = 50
and x1,x2,x3,x4,x5,x6,x7,x8,x9,A1,A2,A3,A4,A5,A6≥0


Iteration-1 Cj 1 2 2 4 3 4 4 1 3 M M M M M M
B CB XB x1 x2 x3 x4 x5 x6 x7 x8 x9 A1 A2 A3 A4 A5 A6 MinRatio
XB/x1
A1 M 100 (1) 1 1 0 0 0 0 0 0 1 0 0 0 0 0 100/1=100
A2 M 50 0 0 0 1 1 1 0 0 0 0 1 0 0 0 0 ---
A3 M 100 0 0 0 0 0 0 1 1 1 0 0 1 0 0 0 ---
A4 M 150 1 0 0 1 0 0 1 0 0 0 0 0 1 0 0 150/1=150
A5 M 50 0 1 0 0 1 0 0 1 0 0 0 0 0 1 0 ---
A6 M 50 0 0 1 0 0 1 0 0 1 0 0 0 0 0 1 ---
Z=500M Zj 2M 2M 2M 2M 2M 2M 2M 2M 2M M M M M M M
Zj-Cj 2M-1↑ 2M-2 2M-2 2M-4 2M-3 2M-4 2M-4 2M-1 2M-3 0 0 0 0 0 0



Positive maximum Zj-Cj is 2M-1 and its column index is 1. So, the entering variable is x1.

Minimum ratio is 100 and its row index is 1. So, the leaving basis variable is A1.

∴ The pivot element is 1.

Entering =x1, Departing =A1, Key Element =1

R1(new)=R1(old)


R2(new)=R2(old)


R3(new)=R3(old)


R4(new)=R4(old) - R1(new)


R5(new)=R5(old)


R6(new)=R6(old)


Iteration-2 Cj 1 2 2 4 3 4 4 1 3 M M M M M
B CB XB x1 x2 x3 x4 x5 x6 x7 x8 x9 A2 A3 A4 A5 A6 MinRatio
XB/x8
x1 1 100 1 1 1 0 0 0 0 0 0 0 0 0 0 0 ---
A2 M 50 0 0 0 1 1 1 0 0 0 1 0 0 0 0 ---
A3 M 100 0 0 0 0 0 0 1 1 1 0 1 0 0 0 100/1=100
A4 M 50 0 -1 -1 1 0 0 1 0 0 0 0 1 0 0 ---
A5 M 50 0 1 0 0 1 0 0 (1) 0 0 0 0 1 0 50/1=50
A6 M 50 0 0 1 0 0 1 0 0 1 0 0 0 0 1 ---
Z=300M+100 Zj 1 1 1 2M 2M 2M 2M 2M 2M M M M M M
Zj-Cj 0 -1 -1 2M-4 2M-3 2M-4 2M-4 2M-1↑ 2M-3 0 0 0 0 0



Positive maximum Zj-Cj is 2M-1 and its column index is 8. So, the entering variable is x8.

Minimum ratio is 50 and its row index is 5. So, the leaving basis variable is A5.

∴ The pivot element is 1.

Entering =x8, Departing =A5, Key Element =1

R5(new)=R5(old)


R1(new)=R1(old)


R2(new)=R2(old)


R3(new)=R3(old) - R5(new)


R4(new)=R4(old)


R6(new)=R6(old)


Iteration-3 Cj 1 2 2 4 3 4 4 1 3 M M M M
B CB XB x1 x2 x3 x4 x5 x6 x7 x8 x9 A2 A3 A4 A6 MinRatio
XB/x9
x1 1 100 1 1 1 0 0 0 0 0 0 0 0 0 0 ---
A2 M 50 0 0 0 1 1 1 0 0 0 1 0 0 0 ---
A3 M 50 0 -1 0 0 -1 0 1 0 1 0 1 0 0 50/1=50
A4 M 50 0 -1 -1 1 0 0 1 0 0 0 0 1 0 ---
x8 1 50 0 1 0 0 1 0 0 1 0 0 0 0 0 ---
A6 M 50 0 0 1 0 0 1 0 0 (1) 0 0 0 1 50/1=50
Z=200M+150 Zj 1 -2M+2 1 2M 1 2M 2M 1 2M M M M M
Zj-Cj 0 -2M -1 2M-4 -2 2M-4 2M-4 0 2M-3↑ 0 0 0 0



Positive maximum Zj-Cj is 2M-3 and its column index is 9. So, the entering variable is x9.

Minimum ratio is 50 and its row index is 6. So, the leaving basis variable is A6.

∴ The pivot element is 1.

Entering =x9, Departing =A6, Key Element =1

R6(new)=R6(old)


R1(new)=R1(old)


R2(new)=R2(old)


R3(new)=R3(old) - R6(new)


R4(new)=R4(old)


R5(new)=R5(old)


Iteration-4 Cj 1 2 2 4 3 4 4 1 3 M M M
B CB XB x1 x2 x3 x4 x5 x6 x7 x8 x9 A2 A3 A4 MinRatio
XB/x4
x1 1 100 1 1 1 0 0 0 0 0 0 0 0 0 ---
A2 M 50 0 0 0 1 1 1 0 0 0 1 0 0 50/1=50
A3 M 0 0 -1 -1 0 -1 -1 1 0 0 0 1 0 ---
A4 M 50 0 -1 -1 (1) 0 0 1 0 0 0 0 1 50/1=50
x8 1 50 0 1 0 0 1 0 0 1 0 0 0 0 ---
x9 3 50 0 0 1 0 0 1 0 0 1 0 0 0 ---
Z=100M+300 Zj 1 -2M+2 -2M+4 2M 1 3 2M 1 3 M M M
Zj-Cj 0 -2M -2M+2 2M-4↑ -2 -1 2M-4 0 0 0 0 0



Positive maximum Zj-Cj is 2M-4 and its column index is 4. So, the entering variable is x4.

Minimum ratio is 50 and its row index is 4. So, the leaving basis variable is A4.

∴ The pivot element is 1.

Entering =x4, Departing =A4, Key Element =1

R4(new)=R4(old)


R1(new)=R1(old)


R2(new)=R2(old) - R4(new)


R3(new)=R3(old)


R5(new)=R5(old)


R6(new)=R6(old)


Iteration-5 Cj 1 2 2 4 3 4 4 1 3 M M
B CB XB x1 x2 x3 x4 x5 x6 x7 x8 x9 A2 A3 MinRatio
x1 1 100 1 1 1 0 0 0 0 0 0 0 0
A2 M 0 0 1 1 0 1 1 -1 0 0 1 0
A3 M 0 0 -1 -1 0 -1 -1 1 0 0 0 1
x4 4 50 0 -1 -1 1 0 0 1 0 0 0 0
x8 1 50 0 1 0 0 1 0 0 1 0 0 0
x9 3 50 0 0 1 0 0 1 0 0 1 0 0
Z=500 Zj 1 -2 0 4 1 3 4 1 3 M M
Zj-Cj 0 -4 -2 0 -2 -1 0 0 0 0 0



Since all Zj-Cj≤0

Hence, optimal solution is arrived with value of variables as :
x1=100,x2=0,x3=0,x4=50,x5=0,x6=0,x7=0,x8=50,x9=50

Min Z=500

also the artificial variable A2 appears in the basis